# Analysis of Hierarchical AoII over unreliable channel: A Stochastic Hybrid System Approach

### v1

@@ IV-B Average AoII under Quadratic Hierarchy
For quadratic hierarchy scheme (21), the average AoII $x_{c}$ of two M/M/1/1 queues over the collision channel at the monitor is summarized as (22b) under certain stability conditions, where the denominator is given in (22c).
The derivation of Corollary 3 appears in Appendix E, where the constants $c_{1}$ and $c_{2}$ are given in (70) as the expected values of AoII at states $1_{m}$ and $2_{m}$, respectively. The associated matrix $\mathbf{D}$ is defined in (71), where the explicit expression of it determinant $\det(\mathbf{D})$ is given in Appendix F and is omitted here for simplicity.
IV-C Sufficient Conditions of Stability over Collision Channel
The stability conditions of the general case (22b) for any $\lambda,\lambda_{c},\mu,\mu_{c}$ are hard to obtain since the stable growth rates $k_{1}$ and $k_{2}$ depend on the relationship between two pairs $(\lambda,\mu)$ and $(\lambda_{c},\mu_{c})$. Therefore, we now present some sufficient conditions for stability control in collision channel and give our finer results for a specially symmetric case where the contender $X^{c}_{t}$ is evenly matched to $X_{t}$, i.e., $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$.
@@ IV-C Sufficient Conditions of Stability over Collision Channel
Assuming that the following conditions are satisfied in the collision channel:
@@
ii). The growth rates $k_{1}$ and $k_{2}$ at the quadratic growth states is limited, i.e., $0<k_{1}<\lambda$ and $0<k_{1}<k_{2}<\mu$,
@@
The proof of Theorem 3 appears in Appendix F. Note that (23) helps to reveal the explicit relationship between $k_{1}$ and $k_{2}$ similar to Theorem 2 with the knowledge of the rates of contender.
@@
We must admit that the whole range of $k_{1}$ and $k_{2}$ ensuring stability might be (much) larger than the intervals given in Theorem 3, which are affected by some unknown contender. However, the explicit relationship between $k_{1}$ and $k_{2}$ given by condition i) are intractable unless we know the relationship of rates between the contender and the source we interest. Nevertheless, it is sufficient to adopt some appropriate constants indicated by the above conditions to achieve two main goals: a). the ability to distinguish different requirements of freshness at different system states; b). the ability to ensure the stability of real-time stochastic communication systems with limited consideration of the effects from environment. In practice, we prefer to choose the parameters $k_{1}$ and $k_{2}$ smaller than $\lambda$ which also are independent of the rates of contender.
@@ V Numerical Results
In this section, we compare the classical results over the M/M/1/1 queue using traditional AoI and AoII metric. Meanwhile, numerical results of the average hierarchical AoII over noisy channel and collision channel are given under different parameters, respectively. The effects of different channel conditions are also analyzed.
Figure 5: Illustration of the comparisons of the average age over the M/M/1/1 queue between AoI and AoII
Figure 6: The average AoII over noisy channel with changing parameters $p$ and $p_{e}$, where the growth rates $k_{1}=1$ and $k_{2}=1.5$. The curves under different utilization factors $\rho=0.2,0.5,0.9$ are given in (a), (b), (c), respectively.
@@ V-A Comparisons between AoI and AoII: The M/M/1/1 queue
1). The M/M/1/1 queue: an update is accepted in the system only if the server is idle, furthermore, new arrivals are discarded when there is an update in service. The average age is given by [5, Equ. (21)], i.e.,
@@
2). The M/M/1/1 queue with preemption: an update-in-service is preempted and discarded if there is a new arrival. The average age is given by [41, Equ. (8)]666also see [43, Equ. 18]., where the source only generates updates of one kind and the service time is set to be exponential, i.e.,
@@
3). The M/M/1/1 queue with abandonment: an update-in-service is abandoned (i.e. discarded without completing service) at rate $\alpha$. The average age is given by [1, Equ. (24)],
@@
We first observe that
@@
The reason of the first inequality is that the AoI metric incorporates a redundant ageing process when the content of update is unchanged (still fresh) at the transmitter. The reason of the second inequality is that the possible change of content ($p<1$) using AoII prolongs the occupancy time of the correct state $0_{c}$ compared to the AoS case ($p=1$). Since the age at $0_{c}$ is always zero, the prolongation helps to further reduce the average.
@@
where the constant $\frac{1}{\mu}$ at the right-hand side can be interpreted as the average age of a generate-at-will queue with Poisson service rate $\mu$. A generate-at-will queue is the queue that a new content-changed update arrives exactly when the current update finishes transmission. Hence, this inequality reveals that the occupancy time of the state $0_{c}$ in (28) in fact brings a reduction of $\frac{1}{\lambda+\mu}$ in the average age.
Moreover, we can further conclude that the average age relationship between (27) and the generate-at-will queue under the same content-change probability $p$, whose average age is $\frac{p}{\mu}$, depends on $p$ and the utilization factor $\rho$.
We illustrate the above results numerically in Fig. 5 by setting the service rate $\mu=1$, where we adopt the natural logarithm of the average age. Note that a larger abandonment rate $\alpha$ leads to a smaller average AoI, whereas a larger content-change probability $p$ gives a larger average AoII.
@@ V-B The effects of noise in the M/M/1/1 queue
2). The range of $p_{e}$ is limited, e.g., $p_{e}<0.4$, $p_{e}<0.6$ and $p_{e}<0.7$ in Fig. 6a, 6b and 6c, respectively, due to the first stability condition given in Theorem 2. Based on this, we could conclude that a poor channel condition may lead to instability even the growth rates are relatively small. In SHSs, it means that any $p_{e}$ violating the stability condition results in a finite escape time of the AoII with non-zero probability.
Furthermore, it is interesting to note that the average AoII at a poor channel condition (for instance, $p_{e}>0.5$) need not always grow monotonically with larger $p$ as seen in Fig. 6c, e.g., the high utilization case. One reasonable guess of this inconsistency could be that the high content-changed probability $p$ leads to relatively small occupancy time in state $0_{e}$ and large occupancy time in state $0_{c}$ when the system state is more likely to transfer into $1_{u}$. Since the difference is not significant between growth rates $k_{1}$ and $k_{2}$, the average age under this circumstance is lower instead.
Figure 7: The average AoII over collision channel in symmetric case, where the growth constants $k_{1}=1$ and $k_{2}=5$ are chosen to satisfy the stability condition in Theorem 3.
V-C The effects of collision in two M/M/1/1 queues
Considering the complexity of (22b), we only illustrate the symmetric case, i.e., $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$, where the stability conditions in Theorem 3 are easy to obtain. The curves of different values of utilization factors (with $\mu=\mu_{c}=10$ fixed) are depicted in Fig. 7 with varying content-changed probability $p$. The average AoII increases monotonically with larger utilization factors as expected since the higher arrival rates of updates from both sources are more likely to cause channel collision.
@@ VI Discussions and Conclusions
In this paper, we consider two different hierarchy schemes, i.e., the linear and quadratic hierarchies, for different discrete states of SHSs, and provide a systematic way to derive the closed-form expressions of average hierarchical AoII. Moreover, we give the stability conditions for the quadratic hierarchy in both channel cases by exploiting the linear systems established by SHSs, which are essential for the existence of the average AoII (also the higher moments).
The derivation of explicit expressions of different age moments in stochastic systems are commonly classified into a set of problems in renewal-reward process. As an alternative of traditional renewal-reward methods, the SHSs method introduced by [9] is considered to be more powerful in that it can not only deal with age-dependent transition rates, which is first studied in [44], also the age-dependent hierarchies that is first considered in our work. However, the limitations of SHSs are also apparent from the collision channel case, i.e., the computation complexity of Theorem 1. To be clearer, one could discuss two simple cases of generalization:
i). variations of multi-level hierarchical AoII with a finer state definitions;
ii). general multi-access cases with $M\geqslant 3$ with a larger finite state space $\mathbb{Q}_{M}$ considering a $k$-way collision.
Both generalizations follow the same analysis procedure, yet the computation cost is enormous due to the order of the associated matrix. Case i) has been discussed after Theorem 2 also in Remark 3. In case ii), the state definition for AoII is more complex than the AoI case in [45]. One should also track the number $k$ of active transmitters as in [45], however, the simplest definition of $\mathbb{Q}_{M}$ is $\mathbb{Q}_{M}=\mathbb{S}_{1}\cup\mathbb{S}_{2}\cup\mathbb{S}_{3}$, where
$\mathbb{S}_{1}$ =$=\{0_{e},0_{c}\},\ \mathbb{S}_{2}=\{M_{m},M_{u}\},$ (34)
$\mathbb{S}_{3}$ =$=\{k_{c},k_{m},k_{u},k_{e}:1\leqslant k\leqslant M-1\}.$
The meanings of states are defined similarly as in the two-node collision case according to their indexes. Hence, the number of states in $\mathbb{Q}_{M}$ is at least $4M$, which in turn gives a set of linear equations with the size of the associated matrix being $3M\times 3M$777the number of irrelevant states is $M$ Although one could solve for the above generalizations with the help of computer software, the expressions could be very complex and no stability conditions could be easily obtained. Based on this, a promising direction on the analysis of continuous AoII in more complex stochastic (queue) systems could be the searches for appropriate bounds.
Stochastic control and policy improvement are not discussed, which are still fresh topics in the AoII literature, especially in the continuous case. Our undergoing works are devoted to address these state-of-the-art problems.
@@ Appendix A Proof of Theorem 1: The Third Part
Since it follows from (7) and Lemma 1 that the change of expected values of $\psi_{\overline{\mathbf{q}}}$ is controlled by an individual stochastic differential equation at each state $\overline{\mathbf{q}}\in\mathbb{Q}$, we only need to rederive the formulas of expected values $\mathbf{v}_{\overline{\mathbf{q}}}$ at quadratic growth states. Without loss of generality, we leverage state $1_{u}$ of the noisy channel case to prove (14).
@@
We can then substitute (39) into Dynkin formula (10), which results in
@@
Assuming that the stability condition is satisfied, the derivative at the left-hand side shall equal to zero as $t\to\infty$. Hence, we obtain
@@
Appendix B Derivations of Corollary 1: The Linear Hierarchy over Noisy Channel
@@
Appendix C Derivations of Corollary 2: The Quadratic Hierarchy over Noisy Channel
By leveraging (13) and (14) at each state, we obtain the set of linear equations of the quadratic case, i.e.,
@@ Appendix C Derivations of Corollary 2: The Quadratic Hierarchy over Noisy Channel
Appendix D Stability Conditions of the Quadratic Hierarchy over Noisy Channel
D-1 Proof of Necessity
We first prove the necessity of these two conditions. Given that the joint state $(\mathbf{q},\mathbf{x})$ is ergodic, the quadratic hierarchy AoII $x_{2}(t)$ converges to a limit $x_{2}$ that can be obtained by (47). Therefore, we are able to establish the necessary conditions from two aspects: 1) the invertibility of matrix $\mathbf{C}$; 2) the positivity of $v^{2}_{\overline{\mathbf{q}}}$ at each state $\overline{\mathbf{q}}$.
Following 1) and 2), we need to calculate the related quantities from (46). The determinant of $\mathbf{C}$ is obtained by the definition

### v2

@@ IV-B Average AoII under Collision Channels
For hierarchy scheme 3, the average AoII $x_{c}$ of two M/M/1/1 queues over the collision channel at the monitor is summarized as (24b) given that the steady-state stability conditions satisfy, where the denominator is given in (24c).
The derivation of Corollary 3 appears in Appendix F, where $c_{1}$ and $c_{2}$ are given in (72) as the expected values of AoII at states $1_{m}$ and $2_{m}$, respectively. The associated matrix $\mathbf{D}$ is defined in (73), where the explicit expression of $\det(\mathbf{D})$ is given in Appendix G and is omitted here for simplicity.
IV-C Sufficient Conditions of Steady-state Stability
For collision case 1), the stability conditions of the general case (24b) for any $\lambda,\lambda_{c},\mu,\mu_{c}$ are hard to obtain since $k_{1}$ and $k_{2}$ depend on the relationship between two pairs $(\lambda,\mu)$ and $(\lambda_{c},\mu_{c})$. Therefore, we now present some sufficient conditions for steady-state stability in the collision case 1) and give our finer results for a symmetric case where the contender $X^{c}_{t}$ is evenly matched to $X_{t}$, i.e., $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$.
@@ IV-C Sufficient Conditions of Steady-state Stability
Assume that the following conditions are satisfied in the collision channel:
@@
ii). The growth constants $k_{1}$ and $k_{2}$ is limited, i.e., $0<k_{1}<\lambda$ and $0<k_{1}<k_{2}<\mu$.
@@
The proof of Theorem 3 appears in Appendix G. Note that (25) helps to reveal the relationship between $k_{1}$ and $k_{2}$ similar to Theorem 2 with the knowledge of the rates of contender.
@@
We recognize that the range of $k_{1}$ and $k_{2}$ ensuring stability might be (much) larger than the intervals given in Theorem 3, which are affected by the contender with unknown parameters. However, the relationship between $k_{1}$ and $k_{2}$ given by condition i) is intractable unless we know the relationship of rates between the contender and the source of interest. Nevertheless, it is sufficient to adopt some appropriate constants given in the above conditions to achieve two main goals: a). The ability to distinguish different requirements of freshness at different system states; b). The ability to ensure the steady-state stability of real-time stochastic communication systems with limited consideration of the effects from environment. In practice, we prefer to choose $k_{1}$ and $k_{2}$ smaller than $\lambda$ which are independent of the rates of contender.
IV-D Discussions on Collision Case 2)
We provide the analysis of collision case 2), where no updates can be successfully decoded after suffering collisions, in the supplementary material of this paper.
@@ V Numerical Results
In this section, we compare the classical results over the M/M/1/1 queue using traditional AoI and AoII metric. Meanwhile, numerical results of the average hierarchical AoII over the noisy and collision channels are given under different parameters, respectively. The effects of different channel conditions are also analyzed.
@@ V-A Comparisons between AoI and AoII: The M/M/1/1 queue
1). The M/M/1/1 queue: an update is accepted in the system only if the server is idle, furthermore, new arrivals are discarded when there is an update in service. The average AoI is given by [2, Equ. (21)], i.e.,
@@
2). The M/M/1/1 queue with preemption: an update-in-service is preempted and discarded if there is a new arrival. The average AoI is given by [41, Equ. (8)]888also see [45, Equ. 18]., where the source only generates updates of one kind and the service time is set to be exponential, i.e.,
@@
3). The M/M/1/1 queue with abandonment: an update-in-service is abandoned (i.e. discarded without completing service) at rate $\alpha$. The average AoI is given by [1, Equ. (24)],
@@
Based on these results, we first observe that
@@
The reason of the first inequality is that the AoI metric incorporates a redundant ageing process when the content of update is unchanged at the transmitter. The reason of the second inequality is that the possible change of content ($p<1$) using AoII prolongs the occupancy time of state $0_{c}$ compared to the AoS case ($p=1$). Since the age at $0_{c}$ is always zero, the prolongation helps to further reduce the average.
@@
where the constant $\frac{1}{\mu}$ can be interpreted as the average age of a generate-at-will queue with Poisson service rate $\mu$. A generate-at-will queue is the queue that a new content-changed update arrives exactly when the current update finishes transmission. Hence, this inequality reveals that the occupancy time of state $0_{c}$ in (30) brings a reduction of $\frac{1}{\lambda+\mu}$ in the average. Moreover, under the same $p$, we can further conclude that the relationship of average age between (29) and the generate-at-will queue depends on $p$ and the utilization factor $\rho$.
Figure 5: Illustration of the comparisons between the average AoI and AoII over the M/M/1/1 queue.
We illustrate the above results numerically in Fig. 5 by setting the service rate $\mu=1$, where we adopt the natural logarithm of the average. Note that a larger abandonment rate $\alpha$ leads to a larger average AoI, and a larger content-change probability $p$ gives a larger average AoII.
@@
Figure 6: The average AoII over noisy channel with changing parameters $p$ and $p_{e}$, where the growth constants $k_{1}=1$ and $k_{2}=1.5$. The curves under different utilization factors $\rho=0.2,0.5,0.9$ are given in (a), (b), (c), respectively.
@@ V-B The effects of noise in the M/M/1/1 queue
2). The range of $p_{e}$ is limited, e.g., $p_{e}<0.4$, $p_{e}<0.6$ and $p_{e}<0.7$ in Fig. 6a, 6b and 6c, respectively, due to the steady-state stability conditions given in Theorem 2. Based on this, we conclude that a poor channel condition may lead to instability even the growth rates are relatively small. In SHSs, it means that any $p_{e}$ violating the stability conditions results in a finite escape time of the AoII with non-zero probability.
Furthermore, when the channel error probability $p_{e}$ is large, e.g., $p_{e}>0.5$, the average AoII does not always monotonically increase in $p$ under the high utilization $\rho=0.9$ as seen in Fig. 6c. One possible reason of this inconsistency is that the occupation time in state $0_{e}$$1_{u}$ for $p=1$ is longer/shorter compared to that of $p=0.8$ or $p=0.6$ under the high $\rho$, which can be derived from (18). Since the growth constant $k_{2}$ being in state $1_{u}$ is larger than the growth constant $k_{1}$ being in state $0_{e}$, the average AoII for $p=1$ is lower instead.
V-C The effects of collisions in two M/M/1/1 queues
Considering the complexity of (24b), we only illustrate the symmetric case, i.e., $\lambda=\lambda_{c}$ and $\mu=\mu_{c}$, where the stability conditions in Theorem 3 are easy to obtain. The curves of different values of utilization factors (with $\mu=\mu_{c}=10$ fixed) are depicted in Fig. 7 with varying content-changed probability $p$. The average AoII increases monotonically with larger utilization factors as expected since the higher arrival rates of updates from both sources are more likely to cause channel collisions. [^9] The reason why the average AoII does not increase monotonically in $p$ under large $\lambda$ is still unclear as of this writing. Nevertheless, we believe that a reasonable explanation still lies in the stationary distributions of discrete state $\mathbf{q}$. To find the exact answer, one could calculate the derivative of $\pi^{\prime}_{2_{u}}$ in (LABEL:collision_distribution) with respect to $p$ to see if a higher $\lambda$ will or not lead to a longer occupancy time in $\pi^{\prime}_{2_{u}}$. However, considering that the explicit expression of $\pi_{2_{u}}$ cannot be easily factorized, it is too complex to obtain the closed-form expression of the derivative.
Figure 7: The average AoII over the collision channel of case 1), where $k_{1}=1$ and $k_{2}=5$ are chosen to satisfy the stability conditions in Theorem 3.
@@ VI Discussions and Conclusions
We first introduce state-dependent AoII processes under different hierarchy schemes. We provide a systematic way to derive the closed-form expressions of the average AoII and obtain explicit results under two different channel models. Moreover, we give the stability analysis based on the SHSs, which are essential for the existence of the average AoII.
As an alternative of renewal-reward methods in obtaining closed-form expressions, the SHSs method introduced by [6] is considered to be more powerful in that it can not only deal with age-dependent transition rates, which is first studied in [35], also the state-dependent ageing processes that are first considered in our work. However, the limitations of SHSs are also apparent as indicated in the collision channel case, i.e., the computation complexity of Theorem 1. To be clear, one can discuss two simple cases of generalization:
i). Variations of multi-level hierarchical AoII processes with a finer state definitions;
ii). General multi-access cases with $M\geqslant 3$ sources under a larger finite state space $\mathbb{Q}_{M}$.
We have discussed i) after Theorem 2, also in Remark 5. For ii), the state definition for the AoII-based multi-access is more complex than the AoI-based multi-access model considered in [44], which in fact makes this generalization intractable. In other words, although both generalizations follow the same analysis procedure in this work, the computation cost for each is enormous due to the order of the matrix associated with the AoII SHSs. To be specific, if we considered a general multi-access case with $M\geqslant 3$ sources, we need to track the number $k$ of sources that have updates transmitting via the channel, which causes a $k$-way collision as noted in [44]. However, by adopting the hierarchical AoII, the simplest definition of discrete states in $\mathbb{Q}_{M}$ is $\mathbb{Q}_{M}=\mathbb{S}_{1}\cup\mathbb{S}_{2}\cup\mathbb{S}_{3}$, where
$\mathbb{S}_{1}$ $=\{0_{e},0_{c}\},\ \mathbb{S}_{2}=\{M_{m},M_{u}\},$ (36)
$\mathbb{S}_{3}$ $=\{k_{c},k_{m},k_{u},k_{e}:1\leqslant k\leqslant M-1\}.$
The meanings of states are defined similarly as in the two-nodes collision model according to their indexes. Hence, the number of states in $\mathbb{Q}_{M}$ is at least $4M$, which in turn gives a set of linear equations with the size of the associated matrix being $3M\times 3M$.101010the number of irrelevant states is $M$. Although we could solve for the above generalization with the help of computer software, the closed-form expressions of average AoII are very complex, and no stability conditions can be easily obtained. A promising direction on the analysis of continuous AoII in such complex stochastic queuing systems would be the searches for appropriate bounds.
@@ Appendix A Proof of Theorem 1: The Third Part
Since it follows from (9) and Lemma 1 that the change of expected values of $\psi_{\overline{\mathbf{q}}}$ is controlled by an individual stochastic differential equation at each state $\overline{\mathbf{q}}\in\mathbb{Q}$, we only need to rederive the formulas of expected values $\mathbf{v}_{\overline{\mathbf{q}}}$ using the linear growth rate function. Without loss of generality, we leverage state $1_{u}$ of the noisy channel case to prove (16).
@@
We can then substitute (41) into Dynkin formula (12), i.e.,
@@
Assuming that the stability condition is satisfied, the derivative shall equal to zero as $t\to\infty$. Hence, we obtain
@@
Appendix B Proof of Positive Recurrence
Let $(\mathbf{q}_{0},\mathbf{x}_{0})$ denotes the initial condition at time $t_{0}$ and $\{t_{i}:i\geqslant 0\}$ be the sequence of transition time. On each interval $t_{i},t_{i+1})$, the AoII at the monitor evolves as
$x_{2}(s)=x_{2}(t_{i})+\int_{t_{i}}^{s}f^{\prime}(\mathbf{q}(r),\mathbf{x}(r),r)\mathrm{d}r,\ \forall s\in[t_{i},t_{i+1}),$ (44)
according to (2). We first discuss Hierarchy Scheme 2. Setting $f^{\prime}=f^{\prime}_{2}$ and taking norms at both sides of (44), we have
[…]
Combining I2) and I3), for $\forall s\in[t_{i},t_{i+1})$ we have
[…]
where $\phi^{x}_{l}$ is the mapping of the continuous state. The inequality follows from the fact that the AoII at the monitor either keeps growing or drops to zero at the transition time $t_{i+1}$.
In the worst case, the second entry of initial condition $\mathbf{x}_{0}=(x_{1}(t_{0}),x_{2}(t_{0}))$ is larger than the constant $\frac{1}{k_{2}}$ and the first term in the maximum function of (50) always dominates as the AoII grows from $i=0$ to i=k1i=k-1, which yields
[…]
Given a positive probability of successful transmission $p_{c}$ in the noisy channel case, the discrete state $\mathbf{q}$ will eventually reach to $0_{c}$ in a finite time with probability one as $t\to\infty$. This condition can be easily verified by calculating the global balance equations for stationary distributions of $\mathbf{q}$. Denoting $t_{k}$ as the last transition time before $\mathbf{q}=0_{c}$, we obtain
[…]
$<\infty,$
[…]
Appendix C Derivations of Corollary 1
@@
Appendix D Derivations of Corollary 2
By leveraging (15) and (16) at each state, we obtain the set of linear equations, i.e.,
@@ Appendix D Derivations of Corollary 2
Appendix E Steady-state Stability Conditions of Hierarchy Scheme 2 over Noisy Channel
E-1 Proof of Necessity
We first prove the necessity of these two conditions. Given that the joint state $(\mathbf{q},\mathbf{x})$ is ergodic, the hierarchical AoII $x_{2}(t)$ converges to a limit $x_{2}$ that can be obtained by (58). Therefore, we are able to establish the necessary conditions from two aspects: 1) the invertibility of matrix $\mathbf{C}$; 2) the positivity of $v^{2}_{\overline{\mathbf{q}}}$ at each state $\overline{\mathbf{q}}$.
Following 1) and 2), we need to calculate the related quantities from (57). The determinant of $\mathbf{C}$ is obtained by